# Choose Your Battles: Distributed Learning Over Multiple Tug of War Games

### v1

@@ 5.2 Distributed Task Allocation with Cooperative Agents
Figure 1: Power control game with (a) $N=50$, (b) $N=4$ players
@@
Figure 2: Performance of Meta-ToP algorithm in multi-channel power control, (a) for a single run, and (b) averaged over 100 realizations
5.3 Activation in Sensor Networks
Let the $N$ sensors be the players where the probability of player $n$ being active is $p_{n}$. The action taken by player $n$ is $x_{n}=1-p_{n}$, i.e., the probability of being off. Clearly, $x_{n}$ is in $[0,1]$. We define the reward function for each player $n$ as
$u_{n}(\boldsymbol{x})=f(P_{n}(\boldsymbol{x}))-\alpha+\beta x_{n}.$
Here $P_{n}(\boldsymbol{x})$ is the probability that player $n$’s observation was successfully transmitted to the destination. This probability depends on the actions of all players as player $n$’s action affects the number of observations it makes, and others’ actions affect how many packets are relayed to the destination. $f(\cdot)$ is the monotonically increasing ‘value’ assigned to observations, which is typically concave since the marginal value of data decreases as the amount of data received increases. Finally, $\beta x_{n}$ is the ‘reward’ for consuming less energy thanks to being off with probability $x_{n}$. This can be thought of as $\beta-\beta p_{n}$, which is a shifted cost due to battery usage. Then $\alpha<0$ is just an offset term that ensure that $u_{n}(\boldsymbol{x})=0$ if $x_{n}=0$.
Unlike in previous applications, here the action $x_{n}$ cannot be thought of as ‘effort’ or the amount of resource used. The reward $u_{n}$ is not necessarily monotone in $x_{n}$ over $[0,1]$. Clearly $P_{n}(\boldsymbol{x})$ decreases and $\beta x_{n}$ increases with increasing $x_{n}$. Hence we can expect a peak at some point if we keep $\boldsymbol{x}_{-n}$ fixed and change $x_{n}$. The aim of the players in this game is to collect as much data as possible but they are limited by their energy consumption. The QoS requirement ensures that sufficient observations is obtained from all sensors. The “best” equilibrium in this example is one where they collect the most data, which is given by the minimal equilibrium.
In this example, the communication of the one-bit ‘out-of-bounds’ signal can easily be done over the existing communication network with little overhead.
6 Simulations
We simulated the games from the previous section. Except for single-run plots, results were averaged over 100 random realizations and plotted along with the interquartile region (the region between the first and the third quartile).
@@ 6 Simulations
Figure 3: Distributed task allocation with $N=100$ agents and $K=10$ tasks for Meta-ToP algorithm with (a) $\rho=0.2,\phi=0.1$, and (b) $\eta(t)=10/(t+1)^{0.6}$
@@
Figure 4: Sensor activation game with $N=10$ sensors for (a) ToP, and (b) FDToP algorithm with stepsize $\eta(t)=1/(t+t_{0})^{\mu}$
6.1 Simulations for Power Control
For the power control game, we first simulate our ToP algorithm for the single-channel case. We randomly generate a diagonal heavy channel gain matrix $C$ where the diagonal elements are uniformly and independently sampled from $[0.2,0.8]$ and the non-diagonal elements are uniformly and independently sampled from $[0,0.2]$. We set $N_{0}=0.1$. Fig. 1(a) compares our algorithms with power control algorithms with $N=50$ players and $\lambda_{n}=0.1$ for each player. We plot the reward of the player with the minimum reward, i.e., the reward of the most dissatisfied player at each time. We assume that the rewards observed are noisy with additive Gaussian measurement noise $\mathcal{N}(0,0.1)$. The upper boundary $B_{n}$ can be set to be large ($B_{n}=1$) so that it has no effect. We use the stepsize sequence $\eta(t)=1/(t+100)$ for our algorithm. The dashed lines represents the QoS requirements.
The algorithms by [1] (Foschini et al. in Fig. 1(a)) and [2] (Zhang et al. in Fig. 1(a)) fail to converge in our case. While [1] cannot handle noise, [2] assumes an unbiased estimator of the inverse SINR, i.e., the inverse reward, which is unavailable to our algorithms. The curve Zhang et al. (2007)++++ denotes the setting in [2], i.e., where Gaussian noise is added to the inverse SINR. As expected, our algorithms converge to the minimal point that satisfies the QoS requirements.
For the multi-channel case (Fig. 2), we simulate the Meta-ToP algorithm for $N=100$ players, $K=10$ channels, and $\lambda_{n}=0.6$ for each player. The rest of the parameters are set as in Fig. 1(a). In Fig. 2(a), we see convergence of the Meta-ToP algorithm for a single run. There are multiple game switches in the beginning leading to frequent returns to zero utility. A feasible game configuration is eventually obtained and the players utilities converge to the QoS requirement.
Fig. 2(b) shows the performance of the Meta-ToP algorithm for three different switching probabilities $\rho,\phi$. The best performance is obtained for $\rho=0.2,\phi=0.1$. Significantly larger probabilities ($\rho=0.6,\phi=0.5$) lead to worse performance since almost random game allocations are being tried after each switch. Smaller probabilities ($\rho=0.02,\phi=0.01$) also have worse performance as this results in very slow exploration. Nevertheless, different $\rho,\phi$ do not drastically change performance, demonstrating the robustness of our algorithm.
@@ 6.2 Simulations for Distributed Task Allocation
For distributed task allocation, we consider $N=100$ agents and $K=10$ tasks. The agents’ efforts are assumed to in $[0,10]$. The constants $\alpha_{g}$ and $\beta_{m,g}$ for the utility function are uniformly and independently sampled from $[1.1,5]$ and $[100,200]$, respectively. The QoS requirement for each player is $0.8$, i.e., $\lambda_{n}=0.8$ for all $n\in\mathcal{N}$. We study the effect of stepsize sequences and the effect of switching probabilities on the performance of the Meta-ToP algorithm (Figure 3).
@@
Fig. 3(b), shows performance of the Meta-ToP algorithm for three different switching probabilities. Unlike in the power control (Fig. 2(b)), performance does not vary significantly with $\rho,\phi$. The performance for $\rho=0.2,\phi=0.1$ is still slightly better, but the convergence rate is similar in all three cases.
@@ 6.3 Simulations for Activation in Sensor Networks
For the sensor activation game ($K=1$), we simulate a network with $N=10$ sensors. Each sensor has a set of multiple routes for sending packets to the destination, and we assume that a transmission is successful if there exists a route where all sensors are active. For each simulation, we randomly generate an Erdős–Rényi graph with edge probability $0.2$ for the sensors. This graph dictates all possible paths from each sensor to its destination. We choose $L=100$ packets and $f(p)=0.8\sqrt{p}$, $\alpha=0.8$ and $\beta=2$. The QoS requirements are chosen as $\lambda_{n}=0.15$ for all players $n\in\mathcal{N}$.
We studied the effects of the step-size sequence on the resulting equilibrium (Fig. 4). In the simulations for the single-channel power control game, the actions never got stuck at the boundary, and the equilibrium is unique.
@@ 7 CONCLUSIONS
We proposed a simple stochastic approximation algorithm that players can use to converge to a point that satisfies their QoS requirements if such a point is feasible. We proved that for Meta-ToW games, our simple algorithm converges to such a desirable point with probability 1. Moreover, for single ToW games, the algorithm converges to the point where the players’ actions are “minimal” with high probability, which is useful when the action represents power or energy. We demonstrate the performance of our algorithm through simulations.
Analyzing ToW games with asynchronous players can remove the dependence on slotted time. It is also an open question what the fastest convergence in ToW games is and whether it is achieved by ToP. Studying better mechanisms of switching between games is another interesting direction.

### v2

@@ 6 Simulations
(c) Percentage of runs converged
Figure 1: Single-channel power control game with (a) $N=50$, (b) $N=4$ players, and (c) varying number of players
@@
(c) Percentage of runs converged
Figure 2: Performance of Meta-ToP algorithm in multi-channel power control
6.1 Simulations for Power Control
For the power control game, we first simulate our ToP algorithm for the single-channel case. We randomly generate a diagonal heavy channel gain matrix $C$ where the diagonal elements are uniformly and independently sampled from $[0.2,0.8]$ and the non-diagonal elements are uniformly and independently sampled from $[0,0.2]$. We set $N_{0}=0.1$. Fig. 1(a) compares our algorithms with power control algorithms with $N=50$ players and $\lambda_{n}=0.1$ for each player. We plot $\min_{n\in\mathcal{N}}u_{n}(\boldsymbol{x}(t))$, i.e., the reward of the player with the minimum reward or the reward of the most dissatisfied player at each time. The rewards observed are noisy with additive Gaussian measurement noise $\mathcal{N}(0,0.1)$. The upper boundary $B_{n}$ is set to be large ($B_{n}=1$) so that it has no effect. We use the stepsize sequence $\eta(t)=1/(t+100)$ for our algorithm. The dashed lines represent the QoS requirements.
The algorithms by [15] (Foschini et al. in Fig. 1(a)) and [40] (Zhang et al. in Fig. 1(a)) fail to converge in our case. While [15] cannot handle noise, [40] assumes an unbiased estimator of the inverse SINR, i.e., the inverse reward, which is unavailable to our algorithms. The curve Zhang et al. (2007)++++ is of the setting in [40], i.e., where Gaussian noise is added to the inverse SINR. The curve Gradient Play corresponds to the algorithm where each player performs gradient ascent on their own utility. Gradient play converges to an equilibrium that does not satisfy the QoS guarantees in this example. As expected, our algorithms converge to the minimal point that satisfies the QoS requirements.
In Fig. 1(c), we plot the rate of convergence for the ToP algorithm for different numbers of players and different QoS requirements. As expected, it takes longer for the players to converge as the number of players and the QoS requirements increase. The first two plots ($N=40,\lambda_{n}=0.1$ and $N=50,\lambda_{n}=0.1$) show that these problems are easier to solve and the players almost immediately converge to an action profile that satisfies the QoS requirements. The remaining QoS requirements are near the boundary of the feasibility region; no action profiles can achieve a utility of $0.16$ for all $N=40$ players, $0.13$ for all $N=50$ players, or $0.11$ for all $N=60$ players; hence, convergence takes longer.
For the multi-channel case (Fig. 2), we simulate the Meta-ToP algorithm for $N=100$ players, $K=10$ channels, and $\lambda_{n}=0.6$ for each player. The rest of the parameters are set as in Fig. 1(a). In Fig. 2(a), we see convergence of the Meta-ToP algorithm for a single run. There are multiple game switches in the beginning, leading to frequent returns to zero utility. A feasible game configuration is eventually obtained and the players’ utilities converge to the QoS requirement.
Fig. 2(b) shows the performance of the Meta-ToP algorithm for three different switching probabilities $\rho,\varphi$. The best performance is obtained for $\rho=0.2,\varphi=0.1$. Significantly larger probabilities ($\rho=0.6,\varphi=0.5$) lead to worse performance since almost random game allocations are being tried after each switch. Smaller probabilities ($\rho=0.02,\varphi=0.01$) also have worse performance as this results in very slow exploration. We study this in more depth in Fig. 2(c), where we plot the rate of convergence for different switching probabilities.
@@ 6.1 Simulations for Power Control
(c) Percentage of runs converged
Figure 3: Distributed task allocation for (a) Meta-ToP with $\rho=0.2,\varphi=0.1$, (b) Meta-ToP with $\eta(t)=10/(t+1)^{0.6}$, and (c) various algorithms
@@
(c) Comparison between algorithms
Figure 4: Performance of various algorithms in sensor activation game with stepsize $\eta(t)=1/(t+t_{0})^{\mu}$
@@ 6.2 Simulations for Distributed Task Allocation
For distributed task allocation, we consider $N=100$ agents and $K=10$ tasks. The agents’ efforts are assumed to be in $[0,10]$. The constants $\alpha_{g}$ and $\beta_{m,g}$ for the utility function are uniformly and independently sampled from $[1.1,5]$ and $[100,200]$, respectively. The QoS requirement for each player is $0.8$, i.e., $\lambda_{n}=0.8$ for all $n\in\mathcal{N}$. We study the effect of stepsize sequences and the effect of switching probabilities on the performance of the Meta-ToP algorithm (Figure 3).
@@
Fig. 3(b), shows performance of the Meta-ToP algorithm for three different switching probabilities. Unlike in the power control (Fig. 2(b)), performance does not vary significantly with $\rho,\varphi$. The performance for $\rho=0.2,\varphi=0.1$ is still slightly better, but the convergence rate is similar in all three cases.
In Fig. 3(c), we compare the rate of convergence of our algorithm with two benchmarks. For our Meta-ToP algorithm, close to 90% of runs converge by $t=25,000$, and 99% of runs converge by $t=100,000$. In one benchmark algorithm, the actions are updated using the ToP algorithm, but the switching happens regularly in fixed intervals. Every $\tau$ timesteps, if not all players have achieved their QoS requirements, a new game configuration is randomly chosen. We tried this algorithm with multiple values of $\tau$ and selected the best-performing ones. While some runs converge for this algorithm, the performance is significantly worse than our Meta-ToP algorithm. As the second benchmark, we compare with an algorithm where actions are updated via gradient play, and switches are made in fixed intervals. This algorithm does not converge to an action profile where all players achieve their QoS requirements.
@@ 6.3 Simulations for Activation in Sensor Networks
For the sensor activation game ($K=1$), we simulate a network with $N=10$ sensors. Each sensor has a set of multiple routes for sending packets to the destination, and we assume that a transmission is successful if there exists a route where all sensors are active. For each simulation, we randomly generate an Erdős–Rényi graph with edge probability $0.2$ for the sensors. This graph dictates all possible paths from each sensor to its destination. We choose $L=100$ packets and $f(p)=0.8\sqrt{p}$, $\alpha=0.8$ and $\beta=0.7$. The QoS requirements are chosen as $\lambda_{n}=0.15$ for all players $n\in\mathcal{N}$.
In the simulations for the single-channel power control game (Subsection VI.A), the actions never got stuck at the boundary, and the equilibrium is unique. Here, since multiple equilibria exist, we studied the effects of the stepsize sequence on the resulting equilibrium (Fig. 4).
@@
In Fig. 4(c), we compare the performance of the ToP algorithm with two gradient-based algorithms. The first algorithm is gradient play, where each player updates their action using gradient ascent on their own utility. The second algorithm is where players update their action using the gradient of the sum of utilities. While our algorithm converges to an action profile where each player achieves their QoS requirements, the gradient-based algorithms converge to a point where at least one player achieves significantly less than their QoS requirement. In both gradient-based algorithms, most players either chose the action $B=1$ or the action 0 at equilibrium.
@@ 7 CONCLUSIONS
We proposed a simple stochastic approximation algorithm that players can use to converge to a point that satisfies their QoS requirements if such a point is feasible. We proved that for Meta-ToW games, our simple algorithm (Meta-ToP) converges to such a desirable point with probability 1. Moreover, for single ToW games, the algorithm converges to the point where the players’ actions are “minimal” with high probability, which is useful when the action represents power or energy. We demonstrate the performance of our algorithm through simulations.
Analyzing ToW games with asynchronous players can remove the dependence on slotted time. It is also an open question what the fastest convergence in ToW games is and whether the ToP achieves it. Studying better mechanisms of switching between games is another interesting direction.